# Functional Percolation: Criticality of Form and Function

### v3

@@ Abstract
Understanding how network structure constrains and enables information processing is a central problem in statistical mechanics of interacting systems [Broadbent1957, Stauffer1994Percolation]. Here we study random networks across the structural percolation transition and analyze how connectivity governs realizable input–output transformations under cascade dynamics. Using Erdős–Rényi networks as a minimal ensemble, we examine structural, functional, and information-theoretic observables as functions of mean degree. We find that the emergence of the giant connected component coincides with a sharp transition in realizable information processing: complex input–output response functions become accessible, functional diversity increases rapidly, output entropy rises, and directed information flow—quantified by transfer entropy—extends beyond local neighborhoods. We term this coincidence of structural, functional, and informational transitions functional percolation, referring to a sharp expansion of the space of realizable input–output functions at the percolation threshold. Near criticality, networks exhibit a Pareto-optimal tradeoff between functional complexity and diversity, suggesting that percolation criticality may provide a general organizing principle of information processing capacity in systems with local interactions and propagating influences.
@@ 1 Introduction
It has been widely recognized that network cascades exhibit hallmarks of universal spreading processes, closely tied to percolation [bak1988self, sethna2001crackling, watts2002simple]. The universality of network cascades appears to arise from the fundamental nature of their minimal physical ingredients: State encoding by system elements, thresholding, information propagation, and energy input to drive the cascade process. These qualities can be found in many neuronal, social, biological, and physical systems. One may also reasonably conjecture that cascades form a general physical substrate for information processing, as node response patterns to inputs constitute Boolean logic functions [Wilkerson2022]. Arguably, due to this dependence on very simple physical constraints and mechanisms, cascades may represent a particularly minimal and broadly applicable class of Boolean network dynamics [Kauffman1969, watts2002simple]. While microscopic interactions are continuous, coarse-graining generically produces state-dependent, threshold-like transitions, making cascade dynamics a widely applicable effective description of information propagation in interacting systems.
As the highest entropy, least structured systems, random networks undergoing cascades provide a minimally-constrained physical basis for information processing, and therefore give an upper bound on structural degrees of freedom available to any realizable system. Thus, studying the emergence of worst-case information processing establishes a broadly applicable baseline of what the heterogeneous systems found in nature can achieve [newman2018networks].
One may also argue for the near-universality of Boolean functions, as they simply characterize patterns of coarse-grained responses to input stimuli, and do not only concern computers, but more fundamentally the simplest units of information [crutchfield1994calculi, shannon1948mathematical]. These functions are also requisite for any system that performs learning or information processing under some level of coarse-graining or decision-making [minsky2017perceptrons, shannon1948mathematical].
@@ 2 The Core Problem: Scale Mismatch Between Form and Function
Cascades in random networks face an additional challenge for information processing. Functions that integrate information from many inputs, having high decision-tree complexity (DTC), are inherently improbable to achieve, as they require connected paths to many input nodes. Such high-DTC functions are exponentially improbable in random networks [Wilkerson2022], despite the fact that they dominate the abstract function space, making them doubly difficult to explore.
@@ 3 Criticality and Functional Percolation
In the universal information processing by cascades in random networks, outlined above, the ability of such systems to perform integrative information processing depends on large-scale connectivity [Wilkerson2022]. Integrative responses are not possible in fragmented random systems. The emergence of global connectivity comes in the form of structural percolation, with the emergence of the giant connected component (GCC) at the critical connectivity $z_{c}$ [Erdos1960RandomGraphs, Stauffer1994Percolation]. Thus the GCC size versus connectivity is a macroscopic order parameter [Stanley1971].
@@
Two entropic order parameters may also be observed to undergo a phase transition at criticality. The output entropy of non-input nodes characterizes the diversity and balance of possible responses — that is, the expressivity of the system under uniform input stimulation. Transfer entropy (TE) characterizes directed information flow through the network, reflecting dynamical responses to inputs in time [shannon1948mathematical, schreiber2000measuring].
@@ 3.1 Model and observables
At criticality, the emergence of the giant connected component is accompanied by a scale-invariant distribution of finite components, which have been shown previously to act as small functional motifs in cascade dynamics [Wilkerson2022]. Due to the simultaneous growth of both DTC and functional diversity near criticality, we observe that this point appears to represent a compromise at Pareto optimality between complexity and diversity of information processing [Langton1990EdgeOfChaos] [Fig. 3]. Exceeding the critical point suppresses exploratory capacity through over-constrained collective dynamics, while connectivity below criticality leads to fragmentation and an inability to realize complex functions. Thus, in the context of the overwhelming size of the Boolean function space, a system optimizes its exploration of diverse, complex functions by operating near criticality.
These considerations motivate the interpretation of criticality as a pivot point between two distinct functional regimes: one supporting the execution of complex, integrative computations, and another enabling efficient exploration of function space. From this perspective, criticality may be viewed as a natural operating point for balancing exploration and exploitation, a distinction that may be framed phenomenologically in terms of reinforcement learning in adaptive or neural systems [Fisher1967, Mora2011, KinouchiCopelli2006].
@@ 4 Discussion and Conclusion
From our results [Figure 1], we observe an agreement between functional and information-theoretic criticality. Transfer entropy can be provides a complementary indicator of directed dynamical influence [schreiber2000measuring]. This gives support to criticality of information processing beyond Boolean functionality. This offers a structural-functional mechanism underlying the neural criticality hypothesis, that optimal information processing happens near criticality, giving us guidance of this behavior from first principles [BeggsPlenz2003, Shew2009DynamicRange, Kinouchi2006DynamicRange]. In the context of the scale mismatch between configuration space and function space discussed above, functional percolation thus appears as a natural way to reconcile limited structural resources with rich information-processing capabilities.
@@
We note that interaction topology and information propagation in networks allow us to begin to connect Boltzmann and Shannon entropies in a direct way, grounded in explicit causal interactions rather than ensemble-level statistical assumptions [sethna2001crackling, Jaynes1957]. At a deeper level, the simultaneous increase of functional complexity, entropy, and information flow at criticality reflects the removal of structural constraints at the percolation transition, which dramatically expands the accessible state space of the system; while this invites a thermodynamic interpretation in terms of effective entropy under coarse-graining, we do not attempt a formal entropy-balance or energetic treatment here. While we are not here making exact formal statements about Landauer bit-erasure or thermodynamics [Landauer1961Irreversibility], it is very compelling to undertake more formal exploration.
We note here that the structure–function scale-mismatch argument above motivates additional search mechanisms by networks beyond static connectivity alone. Criticality itself is likely insufficient to span the exponential scale gap. Dynamics, for example in oscillatory systems, provide a plausible way for systems to reduce dependence on fixed physical connectivity by creating transient virtual computational circuits overlaying the physical topology on short timescales. Kuramoto oscillators, for example, can be said to be ‘connected’—transmitting information—when nearly in phase, but effectively ‘disconnected’ when out of phase [Kuramoto1975, schreiber2000measuring], enabling rapid reconfiguration of functional interactions. One may speculate that such virtual circuit rewiring allows neural systems to explore functional responses using spiking and phase dynamics orders of magnitude faster than physical growth or rewiring [Fries2005-coherence, Varela2001-brainweb, DecoKringelbach2016-metastability]. From a similar perspective, inhibitory neurons may play a complementary role by dynamically reducing effective connectivity, potentially stabilizing networks near criticality in a form of self-organizing criticality (SOC) that facilitates efficient exploration of function space [bak1988self, levina2007dynamical]. While these behaviors have been empirically observed and studied, it is noteworthy that they align closely with mechanisms suggested independently by first-principles arguments and combinatorial constraints.

### v4

@@ Abstract
Understanding how network structure constrains and enables information processing is a central problem in statistical mechanics of interacting systems [3, 26]. Here we study random networks across the structural percolation transition and analyze how connectivity governs realizable input–output transformations under cascade dynamics. Using Erdős–Rényi networks as a minimal ensemble, we examine structural, functional, and information-theoretic observables as functions of mean degree. We find that the emergence of the giant connected component coincides with a sharp transition in realizable information processing: complex input–output response functions become accessible, functional diversity increases rapidly, output entropy rises, and directed information flow—quantified by transfer entropy—extends beyond local neighborhoods. We term this coincidence of structural, functional, and informational transitions functional percolation, referring to a sharp expansion of the space of realizable input–output functions at the percolation threshold. Near criticality, networks exhibit a Pareto-optimal tradeoff between functional complexity and diversity, suggesting that percolation criticality may provide a general organizing principle of information processing capacity in systems with local interactions and propagating influences.
Code and supplementary materials:
https://github.com/galenwilkerson/Functional-Percolation
@@ 1 Introduction
It has been widely recognized that network cascades exhibit hallmarks of universal spreading processes, closely tied to percolation [1, 22, 28]. The universality of network cascades appears to arise from the fundamental nature of their minimal physical ingredients: State encoding by system elements, thresholding, information propagation, and energy input to drive the cascade process. These qualities can be found in many neuronal, social, biological, and physical systems. One may also reasonably conjecture that cascades form a general physical substrate for information processing, as node response patterns to inputs constitute Boolean logic functions [29]. Arguably, due to this dependence on very simple physical constraints and mechanisms, cascades may represent a particularly minimal and broadly applicable class of Boolean network dynamics [11, 28]. While microscopic interactions are continuous, coarse-graining generically produces state-dependent, threshold-like transitions, making cascade dynamics a widely applicable effective description of information propagation in interacting systems.
As the highest entropy, least structured systems, random networks undergoing cascades provide a minimally-constrained physical basis for information processing, and therefore give an upper bound on structural degrees of freedom available to any realizable system. Thus, studying the emergence of worst-case information processing establishes a broadly applicable baseline of what the heterogeneous systems found in nature can achieve [20].
One may also argue for the near-universality of Boolean functions, as they simply characterize patterns of coarse-grained responses to input stimuli, and do not only concern computers, but more fundamentally the simplest units of information [4, 23]. These functions are also requisite for any system that performs learning or information processing under some level of coarse-graining or decision-making [18, 23].
@@
Threshold cascade conditions on networks have been studied extensively, including exact results on seed-size–dependent global cascades [9]. While these works characterize when cascades occur, functional percolation concerns the accessibility and diversity of global response functions enabled by cascades, a distinct question not addressed in prior cascade-percolation analyses
@@ 2 The Core Problem: Scale Mismatch Between Form and Function
Cascades in random networks face an additional challenge for information processing. Functions that integrate information from many inputs, having high decision-tree complexity (DTC), are inherently improbable to achieve, as they require connected paths to many input nodes. Such high-DTC functions are exponentially improbable in random networks [29], despite the fact that they dominate the abstract function space, making them doubly difficult to explore.
@@ 3 Criticality and Functional Percolation
In the universal information processing by cascades in random networks, outlined above, the ability of such systems to perform integrative information processing depends on large-scale connectivity [29]. Integrative responses are not possible in fragmented random systems. The emergence of global connectivity comes in the form of structural percolation, with the emergence of the giant connected component (GCC) at the critical connectivity $z_{c}$ [6, 26]. Thus the GCC size versus connectivity is a macroscopic order parameter [25].
@@
Two entropic order parameters may also be observed to undergo a phase transition at criticality. The output entropy of non-input nodes characterizes the diversity and balance of possible responses — that is, the expressivity of the system under uniform input stimulation. Transfer entropy (TE) characterizes directed information flow through the network, reflecting dynamical responses to inputs in time [23, 21].
@@ 3.1 Model and observables
At criticality, the emergence of the giant connected component is accompanied by a scale-invariant distribution of finite components, which have been shown previously to act as small functional motifs in cascade dynamics [29]. Due to the simultaneous growth of both DTC and functional diversity near criticality, we observe that this point appears to represent a compromise at Pareto optimality between complexity and diversity of information processing [16] [Fig. 3]. Exceeding the critical point suppresses exploratory capacity through over-constrained collective dynamics, while connectivity below criticality leads to fragmentation and an inability to realize complex functions. Thus, in the context of the overwhelming size of the Boolean function space, a system optimizes its exploration of diverse, complex functions by operating near criticality.
These considerations motivate the interpretation of criticality as a pivot point between two distinct functional regimes: one supporting the execution of complex, integrative computations, and another enabling efficient exploration of function space. From this perspective, criticality may be viewed as a natural operating point for balancing exploration and exploitation, a distinction that may be framed phenomenologically in terms of reinforcement learning in adaptive or neural systems [7, 19, 13].
@@ 4 Discussion and Conclusion
From our results [Figure 1], we observe an agreement between functional and information-theoretic criticality. Transfer entropy can provide a complementary indicator of directed dynamical influence [21]. This gives support to criticality of information processing beyond Boolean functionality. It also offers a structural-functional mechanism underlying the neural criticality hypothesis, that optimal information processing happens near criticality, giving us guidance of this behavior from first principles [2, 24, 12]. In the context of the scale mismatch between configuration space and function space discussed above, functional percolation thus appears as a natural way to reconcile limited structural resources with rich information-processing capabilities.
@@
We note that interaction topology and information propagation in networks allow us to begin to connect Boltzmann and Shannon entropies in a direct way, grounded in explicit causal interactions rather than ensemble-level statistical assumptions [22, 10]. At a deeper level, the simultaneous increase of functional complexity, entropy, and information flow at criticality reflects the removal of structural constraints at the percolation transition, which dramatically expands the accessible state space of the system; while this invites a thermodynamic interpretation in terms of effective entropy under coarse-graining, we do not attempt a formal entropy-balance or energetic treatment here. While we are not here making exact formal statements about Landauer bit-erasure or thermodynamics [15], it is very compelling to undertake more formal exploration.
We note here that the structure–function scale-mismatch argument above motivates additional search mechanisms by networks beyond static connectivity alone. Criticality itself is likely insufficient to span the exponential scale gap. Dynamics, for example in oscillatory systems, provide a plausible way for systems to reduce dependence on fixed physical connectivity by creating transient virtual computational circuits overlaying the physical topology on short timescales. Kuramoto oscillators, for example, can be said to be ‘connected’—transmitting information—when nearly in phase, but effectively ‘disconnected’ when out of phase [14, 21], enabling rapid reconfiguration of functional interactions. One may speculate that such virtual circuit rewiring allows neural systems to explore functional responses using spiking and phase dynamics orders of magnitude faster than physical growth or rewiring [8, 27, 5]. From a similar perspective, inhibitory neurons may play a complementary role by dynamically reducing effective connectivity, potentially stabilizing networks near criticality in a form of self-organizing criticality (SOC) that facilitates efficient exploration of function space [1, 17]. While these behaviors have been empirically observed and studied, it is noteworthy that they align closely with mechanisms suggested independently by first-principles arguments and combinatorial constraints.